In part A, you will be accounting for a long-term bond issuance.

Part A – On January 1, 2017, Cheng Inc., issued $200,000 of 8%, 15 year bonds, yielding an effective interest rate of 10%. Semiannual interest is payable on June 30 and December 31. The firm uses effective interest method to amortize any discount or premium.

Required:

1. Determine the issuance price of the bonds and provide the journal entry recorded (impact on the financial statement equation) on January 1, 2017.

2. Prepare the journal entry (impact on the financial statement equation) for the December 31, 2017 interest payment.

3. Determine the interest expense reported on the 2017 income statement.

4. Determine the carrying value of the bonds on the 12/31/2017 balance sheet.
Answer #1
Bond characterstics Amount
1-a) Principal 200,000
interest                8,000
Market interest rate 5%
periods to maturity 30
issue price 169,256
Calculation of bond issue price
Where
i= 5.00%
t= 30
principal * PV of $1 at 5% for 30 yrs =
200,000 * 0.23138        = 46276
interest * PV of ordinary annuity at 5%=
8000 * 15.37245 = 122980
bond issue price 169256
Cash interest increase Carrying
Date paid expense in CV value
1/1/2018 169,256
6/30/2017 8,000 8463 463 169718
12/31/2017 8,000 8486 486 170204
Date Account titles & Explanations Debit Credit
1/1/2017 Cash 169,256
Discount on bonds 30,744
bonds payable 200,000
12/31/2017 interest expense 8486
discount on bonds payable 486
cash 8000
3) interest expense reported - 16949
4) Carrying value of bonds 170204